OpenStudy (anonymous):

ln 6x - ln 2 =5 (logarithm)

OpenStudy (anonymous):

First use the quotient law to obtain: ln(6x/ 2) = 5 or, ln(3x) = 5 or,e^5 = 3x or,x = (e^5)/3

OpenStudy (anonymous):

thanks man appreciate it!